OpenStudy (anonymous):

Complete the square? x^2-7x-6=0

jimthompson5910 (jim_thompson5910):

x^2-7x-6=0 x^2-7x-6+6=0+6 x^2-7x=6 x^2-7x+49/4=6+49/4 ... *** See note below *** (x-7/2)^2 = 73/4 (x-7/2)^2 - 73/4 = 0 Note: on the line marked with 3 stars *** you take half of the x coefficient -7 to get -7/2 and you square it to get 49/4. Then you add this to both sides.

OpenStudy (anonymous):

How would I solve this using the Quadrtic formula?

jimthompson5910 (jim_thompson5910):

Start with \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and plug in a = 1, b = -7, c = -6 to get \[\Large x = \frac{-(-7)\pm\sqrt{(-7)^2-4(1)(-6)}}{2(1)}\] then evaluate \[\Large x = \frac{7\pm\sqrt{49-(-24)}}{2}\] \[\Large x = \frac{7\pm\sqrt{49+24}}{2}\] I'll let you finish.